# A superconvergence result in the RBF-FD method

### v1

@@ 3 The main result
Two different errors will be considered. First, the operator approximation error: For each point we compute the difference between the analytical and RBF-FD $\nabla^{2}$ applied to $u(x,y)$ (Equation 2). For convenience we introduce the vector $\textbf{e}_{\mathrm{op}}$, its $i$-th component being the error at $(x_{i},y_{i})$. We will consider the mean error, which by a slight abuse of notation will be denoted by $||\textbf{e}_{\mathrm{op}}||_{1}$ (despite it actually being the $\ell_{1}$ norm, divided by the number of discretisation nodes $N$). All of the coming results have been observed also in $\ell_{\infty}$ error norm, but will not be displayed for brevity reasons.
@@
Next we consider the solution error, expressing the error of the whole solution procedure when solving our chosen Poisson problem. Analogously to the previous notation we introduce the vector $\textbf{e}_{\mathrm{sol}}$. Solution error convergence can be seen on the right side of Figure 2, which shows our main observation of the paper - convergence orders match those of the operator approximation for odd augmentation degrees $m$, where the error scales as $\propto h^{m-1}$, but we get an order higher for even $m$ approximately $\propto h^{m}$. Additionally, the error spread for even $m$ is noticeably larger, hinting at a different underlying mechanism.
@@
Next, we use this formula to also analyse the behaviour of the solution error. Consider the RBF-FD approximation of $\nabla^{2}$, applied to the solution error: $\nabla^{2}\textbf{e}_{\mathrm{sol}}(i)=\nabla^{2}(\hat{u}(\textbf{x}_{i})-u(\textbf{x}_{i}))=f_{i}-\nabla^{2}u(\textbf{x}_{i})=\textbf{e}_{\mathrm{op}}(i)$. This, together with the fact that the solution error on the boundary is equal to 00 (due to the Dirichlet boundary conditions) implies

### v2

@@ 3 The main result
Two different errors will be considered. First, the operator approximation error: For each point we compute the difference between the analytical and RBF-FD $\nabla^{2}$ applied to $u(x,y)$ (Equation 2). For convenience we introduce the vector $\textbf{e}_{\mathrm{op}}$, its $i$-th component being the error at $(x_{i},y_{i})$. We will consider the mean error, which by a slight abuse of notation will be denoted by $\textbf{e}_{\mathrm{op}}||_{1}$ (despite it actually being the $\ell_{1}$ norm, divided by the number of discretisation nodes $N$). All of the coming results have been observed also in $\ell_{\infty}$ error norm, but will not be displayed for brevity reasons.
@@
Next we consider the solution error, expressing the error of the whole solution procedure when solving our chosen Poisson problem. Analogously to the previous notation we introduce the vector $\textbf{e}_{\mathrm{sol}}$. Solution error convergence can be seen on the right side of Figure 2, which shows our main observation of the paper - convergence orders match those of the operator approximation for odd augmentation degrees $m$, where the error scales as $\propto h^{m-1}$, but we get an order higher for even $m$ - approximately $\propto h^{m}$. Additionally, the error spread for even $m$ is noticeably larger, hinting at a different underlying mechanism.
@@
Next, we use this formula to also analyse the behaviour of the solution error. Consider the RBF-FD approximation of $\nabla^{2}$, applied to the solution error: $\nabla^{2}\textbf{e}_{\mathrm{sol}}(i)=\nabla^{2}(\hat{u}(\textbf{x}_{i})-u(\textbf{x}_{i}))=f_{i}-\nabla^{2}u(\textbf{x}_{i})=\textbf{e}_{\mathrm{op}}(i)$. This, together with the fact that the solution error on the boundary is equal to 0 (due to the Dirichlet boundary conditions) implies